OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the point (6, -3) and has a slope of 1/2 y - 6 = 1/2 (x + 3) y = 1/2 x - 6 y+ 3 = 1/2 (x - 6) x - 2y = 12

OpenStudy (anonymous):

they all seem to have a slope of 1/2 so you might try plugging in (6-3) into all of 'em and see which one holds true...

OpenStudy (anonymous):

i did but i got confused .

OpenStudy (anonymous):

for example, 6 is your x coordinate and -3 is your y coordinate. plug it in to the first equation: y - 6 =1/2(x+3) (-3) -6 = 1/2(6+3) is this a true statement? simplify if you need to.

OpenStudy (anonymous):

hello?